OpenStudy (jazzyluvsu123):

ok I have got my answer for this problem but I don't know what I need to do after @jtvatsim 5&2/3 x 8&1/4= 187/4 <--- what do I do after this do I divide or leave it like that ??? @satellite73 and @timaashorty

OpenStudy (anonymous):

if you start with a mixed number, you probably want to turn it back in to a mixed number

OpenStudy (jazzyluvsu123):

so divide right!

OpenStudy (anonymous):

yes

OpenStudy (jazzyluvsu123):

ok stay so I can check my answer with u

OpenStudy (anonymous):

i think your answer is not right r you could have said \[5+8=13\] and \[\frac{2}{3}+\frac{1}{4}=\frac{11}{12}\]

OpenStudy (anonymous):

the common denominator for \(3\) and \(4\) is \(12\) not \(4\) so there is a mistake somewhere

OpenStudy (jazzyluvsu123):

well my teacher makes us like turn fractions into this 17/1 x 11/4=187/4

OpenStudy (anonymous):

answer should be \[13\tfrac{11}{12}\]

OpenStudy (timaashorty):

as an improper fraction it's right...

OpenStudy (anonymous):

ooohh my fault, it is times, not plus sorry

OpenStudy (jazzyluvsu123):

np

OpenStudy (anonymous):

you are right, my mistake

OpenStudy (jazzyluvsu123):

so I leave my answer like it is ? or no

OpenStudy (anonymous):

now you can turn it back to a mixed number, which is probably what your teacher wants

OpenStudy (jazzyluvsu123):

how ??

OpenStudy (timaashorty):

like you said before by dividing

OpenStudy (anonymous):

like you said, divide

OpenStudy (timaashorty):

\[187 \div 4 = ?\]

OpenStudy (jazzyluvsu123):

igot 46.75

OpenStudy (timaashorty):

you change the decimal into a fraction

OpenStudy (anonymous):

leave the remainder as a fraction you get 46 and a remainder of 3, so \(46\tfrac{3}{4}\)

OpenStudy (jazzyluvsu123):

ok thanks I will do my other answer and then post them on here to see if they are right :) thanks u guys

OpenStudy (timaashorty):

All thanks go to satellite(:

OpenStudy (anonymous):

hardly

OpenStudy (jazzyluvsu123):

umm.. for one of my answers I got 4.8 what would the fraction version be ???

OpenStudy (jazzyluvsu123):

@satellite73 or @timaashorty